# Index Terms—  MRI, implicit data crimes, parallel imaging, reproducing kernel Hilbert spaces

### v1

@@ Abstract
Magnetic Resonance Imaging (MRI) diagnoses and manages a wide range of diseases, yet long scan times drive high costs and limit accessibility. AI methods have demonstrated substantial potential for reducing scan times, but despite rapid progress, clinical translation of AI often fails. One particular class of failure modes, referred to as implicit data crimes, are a result of hidden biases introduced when MRI datasets incompletely model the MRI physics of the acquisition. Previous work identified data crimes resulting from algorithmic completion of k-space with parallel imaging and drew on simulation to demonstrate the resulting downstream biases. This work proposes a mathematical framework to re-characterize the problem as one of error reduction during interpolation between sets of evaluation coordinates. We establish a generalized matrix-based definition of the reconstruction error upper bound as a function of the input sampling pattern. Experiments on relevant sampling pattern structures demonstrate the relevance of the framework and suggest future directions for analysis of data crimes.
@@ 2 Theory
2.1 RKHS-based MRI Reconstruction
@@
2.2 Experiment Error Bound
Consider a linear retrospective experiment as illustrated in Figure 3. Reference data are prospectively acquired at points $S_{\mathrm{pro}}$ then reconstructed to a fully-sampled grid $S_{\mathrm{all}}$. Test reconstruction does the same after first interpolating to an intermediate pattern $S_{\mathrm{retro}}$. Observe that this chain of reconstructions is equivalent to direct interpolation between the subsets.
@@ 2.2 Experiment Error Bound
$=\left|\sum_{{\mathbf{z}}_{r}}\sum_{{\mathbf{z}}_{p}\in S_{\mathrm{pro}}}f({\mathbf{z}}_{p})w({\mathbf{z}}_{p},{\mathbf{z}}_{r})w({\mathbf{z}}_{r},{\mathbf{z}})-\sum_{{\mathbf{z}}_{p}\in S_{\mathrm{pro}}}f({\mathbf{z}}_{p})w({\mathbf{z}}_{p},{\mathbf{z}})\right|$ (7)
@@
When $S_{\mathrm{pro}}=S_{\mathrm{all}}$, the power function describes the single-stage interpolation error of the retrospective experiment. Such a setting is “data-crime free” as all k-space points are Nyquist-rate samples of the data. On the other hand, when $S_{\mathrm{pro}}=S_{\mathrm{retro}}$, the retrospective experiment is trivial, requiring no new interpolation. This case maximizes the “data-crime”, potentially (aside from possible regularization effects) suggesting an error-free result.
@@ 3 Methods
Using the BART software toolbox [bart], a 4-coil, $\times32$ phantom was created in k-space so as to avoid an inverse crime [inverse_crime_mri]. We add noise to the data for an overall SNR of 35dB. Regularization set to $10^{-6}$ times the average eigenvalue of $M(S_{\mathrm{all}},S_{\mathrm{all}})$ was added along the diagonal of all inverted matrices for all experiments. We apply the computational framework in Figure 3 to compare a few pattern combinations: a fully-sampled prospective grid versus an $R_{\mathrm{pro}}=2$ 2D uniform sampling pattern for an $R_{\mathrm{retro}}=3$ uniform 2D pattern; and an $R_{\mathrm{pro}}=4$ CAIPIRINHA [breuer_controlled_2005] 3D prospective pattern versus an $R_{\mathrm{pro}}=4$ random 3D pattern for an $R_{\mathrm{retro}}=4$ random 3D pattern.
@@
Fig. 4: Demonstration of direct interpolation of retrospective experiment error and power function calculation for 2D sampling patterns. Our analysis detects subtle changes in the image-domain error, as well as the magnitudes of the experiment power function when introducing this prospective pattern.
@@ 4 Results
Figure 4 demonstrates the result of introducing a prospective 2D sampling pattern to a retrospective sampling experiment. An acceleration of $R_{\mathrm{pro}}=2$ is a standard clinical practice, and in fact, parallel imaging reconstructs missing k-space points effectively for a uniformly-spaced pattern. Still, the experiment power function, measuring interpolation difficulty between this pattern and an $R_{\mathrm{retro}}=3$ uniform pattern along the same dimension, subtly but perceptibly is suppressed in magnitude compared to using a fully-sampled prospective dataset.
@@
Figure 5 demonstrates combinations of 3D Cartesian sampling patterns, which might vary across two spatial dimensions. Interpolation to a random 3D pattern from a CAIPIRINHA grid generates a power function that is structured according to interpolation gaps in k-space. As this pattern minimizes gaps between k-space samples in 2D, it leads to a stable prospective reconstruction, such that the power function resembles the structure of the 3D pattern. A final example interpolates between two distinct random 3D patterns, outputting a power function that anticipates k-space error and contains structural features of both patterns.
5 Discussion
@@
Experiments visualize the connection between the power function derived above and the choice of sampling pattern pairs in the PI inverse crime. RKHS weights and their corresponding power functions are expensive to compute, requiring the solution to a linear system whose entries are $\Theta(N^{2})$ in each dimension of the input data, and requiring up to $O(N^{6})$ time to solve the system for an $N\times N$ image and even longer for 3D data. For this phantom case, however, RKHS theory successfully predicts the structure of retrospective experiment error for standard MR k-space sampling approaches.
@@
Composite power functions give rigor to the related data crime phenomenon exhibited in [frenklak2024parallel]. The analysis is general and could be applied to arbitrary (e.g., non-Cartesian) sampling patterns. Upper bounds give useful information about the distribution of error in k-space, possibly aiding the design of future reconstruction regularizers and sampling pattern optimizers, as well as predicting the degree of bias due to use of sub-sampled data. While the high computational cost of RKHS interpolaters limits the immediate application of this work, there are clear connections to low-dimensional reconstructors such as GRAPPA and image-domain ones such as SENSE [Deshmane2012Parallel]. Subsequent analysis should attempt to bridge linear error analysis both to those methods and to non-linear AI-based reconstruction, helping researchers to improve the rigor of their algorithmic pipelines.
6 Conclusion
In conclusion, we presented a theoretical framework for analyzing implicit data crimes in multi-coil MRI arising from using PI-completed data for retrospective sub-sampling experiments. Our theory, based on RKHS, upper-bounds the error in linear reconstructors and could be used to detect and assess the degree of bias due to this implicit data crime.

### v2

# Theoretical Bounds on Parallel Imaging Implicit Data Crimes in an MRI Reproducing Kernel Hilbert Space

@@ Abstract
Magnetic Resonance Imaging (MRI) diagnoses and manages a wide range of diseases, yet long scan times drive high costs and limit accessibility. AI methods have demonstrated substantial potential for reducing scan times, but despite rapid progress, clinical translation of AI often fails. One particular failure mode, referred to as implicit data crimes, is a result of hidden biases introduced when MRI datasets incompletely model the MRI physics of the acquisition. Previous work identified data crimes resulting from algorithmic completion of k-space with parallel imaging and drew on simulation to demonstrate the resulting downstream biases. This work proposes a mathematical framework to re-characterize the problem as one of error reduction during interpolation between sets of evaluation coordinates. We establish a generalized matrix-based definition of the reconstruction error upper bound as a function of the input sampling pattern. Experiments on relevant sampling pattern structures demonstrate the relevance of the framework and suggest future directions for analysis of data crimes.
footnotetext: © 2026 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ 2 Theory
2.1 Background: RKHS-based MRI Reconstruction
@@
2.2 Error Bound Extension to Data-Crime
Consider a linear retrospective experiment as illustrated in Figure 3. Noisy reference data are prospectively acquired at points $S_{\mathrm{pro}}$ then reconstructed to a fully-sampled grid $S_{\mathrm{all}}$. Test reconstruction does the same after first interpolating to an intermediate pattern $S_{\mathrm{retro}}$. Observe that this chain of reconstructions is equivalent to direct interpolation between the subsets.
@@ 2.2 Error Bound Extension to Data-Crime
$\begin{split}{}&=\left|\sum_{{\mathbf{z}}_{r}}\sum_{{\mathbf{z}}_{p}\in S_{\mathrm{pro}}}f({\mathbf{z}}_{p})w({\mathbf{z}}_{p},{\mathbf{z}}_{r})w({\mathbf{z}}_{r},{\mathbf{z}})\right.\\ &\quad\quad\quad{}-\left.\sum_{{\mathbf{z}}_{p}\in S_{\mathrm{pro}}}f({\mathbf{z}}_{p})w({\mathbf{z}}_{p},{\mathbf{z}})\right|\end{split}$ (7)
@@
When $S_{\mathrm{pro}}=S_{\mathrm{all}}$, the power function describes the single-stage interpolation error of the retrospective experiment. Such a setting is “data-crime free” as all k-space points are Nyquist-samples of the data. On the other hand, when $S_{\mathrm{pro}}=S_{\mathrm{retro}}$, the retrospective experiment is trivial, requiring no new interpolation. This case maximizes the “data-crime”, potentially (aside from possible regularization effects) suggesting an error-free result.
The norm of columns of the $\Delta W$ matrix gives the noise amplification term $\nu_{j_{b}}(\mathbf{x}_{b})=\|\Delta w_{b}(S_{A},S_{B})\|_{2}$. It is a vector magnitude of difference in weights for output coordinate $(j_{b},\mathbf{x}_{b})$, characterizing the stability of interpolating experiment error there.
@@ 3 Methods
Using the BART software toolbox [bart], a 4-coil, $\times32$ phantom was created in k-space so as to avoid an inverse crime [inverse_crime_mri]. We add noise to the data for an overall SNR of 35dB. Regularization set to $10^{-5}$ times the average eigenvalue of $M(S_{\mathrm{all}},S_{\mathrm{all}})$ was added along the diagonal of all inverted matrices for all experiments. We apply the computational framework in Figure 3 to compare a few pattern combinations: a fully-sampled prospective grid versus an $R_{\mathrm{pro}}=2$ 2D uniform sampling pattern for an $R_{\mathrm{retro}}=3$ uniform 2D pattern; and an $R_{\mathrm{pro}}=4$ CAIPIRINHA [breuer_controlled_2005] 3D prospective pattern versus an $R_{\mathrm{pro}}=4$ random 3D pattern for an $R_{\mathrm{retro}}=4$ random 3D pattern.
@@
Fig. 4: Experiment error and power functions for 2D sampling patterns. Our data-independent power function predicts k-space structure of experiment error, and its magnitude is suppressed due to the introduction of a prospective pattern.
@@ 4 Results
Figure 4 demonstrates the result of introducing a prospective 2D sampling pattern to a retrospective sampling experiment. An acceleration of $R_{\mathrm{pro}}=2$ is a standard clinical practice, and in fact, parallel imaging reconstructs missing k-space points effectively for a uniformly-spaced pattern. Still, the experiment power function, measuring interpolation difficulty between this pattern and an $R_{\mathrm{retro}}=3$ uniform pattern along the same dimension, subtly but perceptibly is suppressed in magnitude compared to using a fully-sampled prospective dataset. This follows from the reduced dimension of all functions interpolated from the prospective pattern, that is, a subspace projection induced by that pattern.
@@
Figure 5 demonstrates combinations of 3D Cartesian sampling patterns, which might vary across two spatial dimensions. Interpolation to a random 3D pattern from a CAIPIRINHA grid generates a power function that is structured according to interpolation gaps in k-space. As this pattern minimizes gaps between k-space samples in 2D, it leads to a stable prospective reconstruction, such that the power function resembles the structure of the 3D pattern. A final example interpolates from a less stable 3D pattern, outputting a power function with some regions of reduced magnitude, anticipating subtle changes in experimental error.
5 Discussion and Conclusion
@@
Experiments visualize the connection between the power function derived above and the choice of sampling pattern pairs in the PI inverse crime. RKHS weights and their corresponding power functions are expensive to compute, requiring the solution to a linear system whose entries are $\Theta(N^{2})$ in each dimension of the input data, and requiring up to $O(N^{6})$ time to solve the system for an $N\times N$ image and even longer for 3D data. Matrix-free programs may help make these solutions tractable for full-size problems. For this phantom case, however, RKHS theory successfully predicts the structure of retrospective experiment error for standard MR k-space sampling approaches.
@@
Composite power functions give rigor to the related data crime phenomenon exhibited in [frenklak2024parallel]. The analysis is general and could be applied to arbitrary (e.g., non-Cartesian) sampling patterns. Upper bounds give useful information about the structure of error in k-space, possibly inspiring future reconstruction regularizers and sampling pattern optimizers, as well as mitigation tools that predict bias due to the PI data crime. While the high computational cost of RKHS interpolaters limits immediate application, these methods relate closely to low-dimensional reconstructors such as GRAPPA and image-domain ones such as SENSE [Deshmane2012Parallel]. Subsequent analysis should bridge linear error analysis to those methods and to non-linear AI-based reconstruction, helping researchers to improve the rigor of their algorithmic pipelines.